### v1

# Complementary polynomials in quantum signal processing

Abstract
Quantum signal processing is a framework for implementing polynomial functions on quantum computers. To implement a given polynomial $P$, one must first construct a corresponding complementary polynomial $Q$. Existing approaches to this problem employ numerical methods that are not amenable to explicit error analysis. We present a new approach to complementary polynomials using complex analysis. Our main mathematical result is a contour integral representation for a canonical complementary polynomial. The integral representation on the unit circle has a particularly simple and efficacious Fourier analytic interpretation, which we use to develop a Fast Fourier Transform-based algorithm for the efficient calculation of $Q$ in the monomial basis with explicit error guarantees. Numerical evidence that our algorithm outperforms the state-of-the-art optimization-based method for computing complementary polynomials is provided.
1 Introduction
Emails: bjorn.berntson@riverlane.com, christoph.sunderhauf@riverlane.com
Quantum signal processing (QSP) [1] and its extensions [2, 3] describe a simple single-qubit parameterised circuit that applies a chosen polynomial to a scalar. They have become indispensable in state-of-the-art quantum algorithms because the circuits can be lifted from scalars to arbitrary matrices, called quantum singular value transformation (QSVT) [4, 5]. The matrix is embedded inside a larger unitary, called a block encoding. The polynomial is then applied to all singular values of the matrix simultaneously, which can lead to a quantum speedup compared to classical evaluation of matrix functions. QSVT has had tremendous impact in quantum computing since polynomials may be used to approximate a wide variety of functions. Thereby, this family of algorithms encompasses many prior quantum algorithms [6], including those for Hamiltonian simulation [7], solving linear systems [8], phase estimation [9], and amplitude amplification [10], often even improving the prior algorithm.
There are different parameterizations of QSP [1, 4, 11]. The fundamental idea underlying each is that certain polynomials may be realized within a matrix element of a finite product of single-qubit unitaries [1, 3, 12]. Determining the parameters of the unitaries required to implement a particular polynomial has proven to be a challenging problem for which several numerical methods have been developed [11, 13, 14, 15, 16]. The generalized quantum signal processing (GQSP) framework [^1] For future convenience, in particular to relate (1.2) to other QSP conventions, we have written (1.2) as the transpose of [3, Eq. 7]. of Motlagh and Wiebe [3] has, as we describe below, clarified the mathematical structure of this factorization problem.
Theorem 1 (Generalized quantum signal processing, [3]).
Let $P\in\mathbb{C}[z]$ such that $\deg P=d\in\mathbb{Z}_{\geq 1}$ and $\lvert P(z)\rvert\leq 1$ on $\mathbb{T}\coloneqq\{z\in\mathbb{C}:\lvert z\rvert=1\}$. Then, there exists $Q\in\mathbb{C}[z]$ such that $\deg Q=d$ and
$\lvert P(z)\rvert^{2}+\lvert Q(z)\rvert^{2}=1\quad(z\in\mathbb{T})$ (1.1)
holds. Moreover, there exist parameters $\lambda\in(-\pi,\pi]$ and $(\theta_{j})_{j=0}^{d},(\phi_{j})_{j=0}^{d}\in(-\pi,\pi]^{d+1}$ such that
$\left(\begin{array}[]{cc}P(z)&Q(z)\\ &*\end{array}\right)=\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}(\lambda+\phi_{0})}\cos\theta_{0}&\mathrm{e}^{\mathrm{i}\lambda}\sin\theta_{0}\\ \mathrm{e}^{\mathrm{i}\phi_{0}}\sin\theta_{0}&-\cos\theta_{0}\end{array}\right)\Bigg{[}\prod_{j=1}^{d}\left(\begin{array}[]{cc}z&0\\ 0&1\end{array}\right)\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}\phi_{j}}\cos\theta_{j}&\sin\theta_{j}\\ \mathrm{e}^{\mathrm{i}\phi_{j}}\sin\theta_{j}&-\cos\theta_{j}\end{array}\right)\Bigg{]}\quad(z\in\mathbb{T}),$ (1.2)
where * indicates the precise form of the matrix elements is immaterial, holds.
In (1.2), we call $Q$ a complementary polynomial to $P$ and the parameters $\lambda$, $(\phi_{j})_{j=0}^{d}$, $(\theta_{j})_{j=0}^{d}$ phase factors. It is shown in [3] that once a complementary polynomial is known, it is mathematically trivial to construct the phase factors to realize (1.2). This remarkable fact stands in contrast to the paradigm of that for standard QSP where, with or without knowledge of the complementary polynomial, one must recourse to numerical methods to determine the phase factors. GQSP subsumes the Laurent [11] and real formulations of QSP, as we show in Appendix A; see also [7]. Thus, the conceptual advantage of GQSP, that the challenge in realizing any polynomial lies wholly in computing a complementary polynomial, descends to standard QSP. Further, the phase factors required in lifted algorithms operating on matrices like QSVT follow immediately from those in QSP. By the above discussion, the fundamental obstacle to representing polynomials in both the GQSP and standard QSP as well as QSVT frameworks is the following problem.
Problem 1 (Complementary polynomials problem).
Given $P\in\mathbb{C}[z]$ satisfying the conditions of Theorem 1, find $Q\in\mathbb{C}[z]$ in the monomial basis, such that $\deg Q=\deg P$ and (1.1) holds.
Previous approaches to this problem rely on root-finding [4], Prony’s method [15], or optimization [3]. In contrast, we construct the first exact representation of a canonical complementary polynomial $Q$, valid throughout the entire complex plane, in the form of a set of contour integrals. The contour integral representation of $Q$ on $\mathbb{T}$ can be rephrased in the language of Fourier analysis. We combine this Fourier analytic interpretation and the fast Fourier transform to develop efficient numerical algorithms to compute $Q$ in the monomial basis.
Exact and explicit error analysis of our algorithms is performed, providing a rigorous upper bound on the classical runtime as a linear function, up to logarithmic factors, of the error and degree. Numerical results from our reference implementation further demonstrate the practicality and competitiveness of our algorithm. We emphasize that existing numerical approaches to the construction of complementary polynomials, which we describe in Section 1.3 below, rely on heuristics and so are not amenable to such error analysis.
In the remainder of this introduction, we state our principal results, give remarks on their proofs, describe related literature, introduce notation used in the main text, and outline the structure of the paper.
1.1 Statement of results
We first construct a set of contour integral representations for $Q$. Let
$P(z)=\sum_{n=0}^{d}p_{n}z^{n}\quad(p_{0}\neq 0);$ (1.3)
the restriction that $p_{0}\neq 0$ is imposed without loss of generality as $\lvert z^{n}P(z)\rvert=\lvert P(z)\rvert$ holds for all $z\in\mathbb{T}$, $n\in\mathbb{Z}$. Our principal result is the following theorem, giving representations of $Q$ on $\mathbb{D}\coloneqq\{z\in\mathbb{C}:\lvert z\rvert<1\}$, $\mathbb{T}$, and $\mathbb{C}\setminus\overline{\mathbb{D}}$.
Theorem 2 (Contour integral representation of the canonical complementary polynomial).
Suppose $P$ satisfying the assumptions of Problem 1 is given in the form (1.3). Let $d_{0}\in\mathbb{Z}_{\geq 0}$ be the number of roots of $1-\lvert P(z)\rvert^{2}$ on $\mathbb{T}$, not counting multiplicity, and $\{(t_{j},2\alpha_{j})\}_{j=1}^{d_{0}}$ be the corresponding roots and multiplicities, which are necessarily even. Then,
$Q(z)=$ ${Q}_{0}(z)\exp\Bigg{(}\frac{1}{4\pi\mathrm{i}}\int_{\mathbb{T}}\frac{z^{\prime}+z}{z^{\prime}-z}\log\bigg{(}\frac{1-\lvert P(z^{\prime})\rvert^{2}}{\lvert{Q}_{0}(z^{\prime})\rvert^{2}}\bigg{)}\frac{\mathrm{d}z^{\prime}}{z^{\prime}}\Bigg{)}$ $z\in\mathbb{D}$ (1.4a)
$Q(z)=$ ${Q}_{0}(z)\exp\Bigg{(}\frac{1}{4\pi\mathrm{i}}\,\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{T}}\frac{z^{\prime}+z}{z^{\prime}-z}\log\bigg{(}\frac{1-\lvert P(z^{\prime})\rvert^{2}}{\lvert{Q}_{0}(z^{\prime})\rvert^{2}}\bigg{)}\frac{\mathrm{d}z^{\prime}}{z^{\prime}}+\frac{1}{2}\log\bigg{(}\frac{1-\lvert P(z)\rvert^{2}}{\lvert{Q}_{0}(z)\rvert^{2}}\bigg{)}\Bigg{)}$ $z\in\mathbb{T}$ (1.4b)
$Q(z)=$ $\frac{1-P(z)P^{*}(1/z)}{Q_{0}^{*}(1/z)}\exp\Bigg{(}\frac{1}{4\pi\mathrm{i}}\int_{\mathbb{T}}\frac{z^{\prime}+z}{z^{\prime}-z}\log\bigg{(}\frac{1-\lvert P(z^{\prime})\rvert^{2}}{\lvert{Q}_{0}(z^{\prime})\rvert^{2}}\bigg{)}\frac{\mathrm{d}z^{\prime}}{z^{\prime}}\Bigg{)}$ $z\in\mathbb{C}\setminus\overline{\mathbb{D}}$, (1.4c)
where
${Q}_{0}(z)\coloneqq\prod_{j=1}^{d_{0}}(z-t_{j})^{\alpha_{j}},$ (1.5)
the integration contour $\mathbb{T}$ is positively-oriented, and the dashed integral indicates a principal value prescription with respect to the singularity $z^{\prime}=z$ on $\mathbb{T}$, solves Problem 1. Moreover, (1.4c) is, up to a multiplicative phase, the unique solution of Problem 1 with no roots in $\mathbb{D}$.
Theorem 2 provides an exact representation of $Q$ in Problem 1, in canonical form: all roots lie outside of $\mathbb{D}$. In the important special case where $P$ has real coefficients, a canonical complementary polynomial with real coefficients exists. The precise statement, a mild generalization of a result in [13, Section 8], is now given.
Corollary 2.1 (Real complementary polynomials).
Let $P\in\mathbb{R}[z]$ satisfying the conditions of Problem 1 be given. Then, the canonical complementary polynomial fulfills $Q\in\mathbb{R}[z]$, up to a multiplicative phase.
To obtain $Q$ explicitly in the monomial basis, i.e., to solve Problem 1, it suffices to evaluate (1.4c) at any $d+1$ distinct points of $\mathbb{C}$ and employ the Lagrange interpolation formula. Our numerical approach is based on interpolation through roots of unity; this is equivalent to a discrete Fourier transform. The following corollary of Theorem 2 establishes a Fourier analytic variant of the integral representation (1.4b) of $Q$ on $\mathbb{T}$, which we will later use to evaluate $Q$ at the roots of unity.
Corollary 2.2 (Fourier analytic variant of Theorem 2 on $\mathbb{T}$).
The representation (1.4b) of $Q$ on $\mathbb{T}$ is equivalent to
$Q(\mathrm{e}^{\mathrm{i}\theta})={Q}_{0}(\mathrm{e}^{\mathrm{i}\theta})\exp\Bigg{(}\Pi\bigg{[}\log\bigg{(}\frac{1-\lvert P(\mathrm{e}^{\mathrm{i}\theta})\rvert^{2}}{\lvert{Q}_{0}(\mathrm{e}^{\mathrm{i}\theta})\rvert^{2}}\bigg{)}\bigg{]}\Bigg{)}\quad(\theta\in(-\pi,\pi]),$ (1.6)
where $\Pi$ is the Fourier multiplier defined by
$\Pi[\mathrm{e}^{\mathrm{i}n\theta}]\coloneqq\begin{cases}\mathrm{e}^{\mathrm{i}n\theta}&n\in\mathbb{Z}_{>0}\\ \frac{1}{2}&n=0\\ 0&n\in\mathbb{Z}_{<0}.\end{cases}$ (1.7)
Due to (1.7), we have
$\Pi\Bigg{[}\sum_{n\in\mathbb{Z}}a_{n}\mathrm{e}^{\mathrm{i}n\theta}\Bigg{]}=\frac{1}{2}a_{0}+\sum_{n=1}^{\infty}a_{n}\mathrm{e}^{\mathrm{i}n\theta},$ (1.8)
and Corollary 2.2 shows that, essentially, constructing $Q$ on $\mathbb{T}$ consists in evaluating the Fourier coefficients of the function $\log\big{(}\frac{1-\lvert P(\mathrm{e}^{\mathrm{i}\theta})\rvert^{2}}{\lvert{Q}_{0}(\mathrm{e}^{\mathrm{i}\theta})\rvert^{2}}\big{)}$.
Numerical methods
Corollary 2.2 suggests a practical numerical method to compute an approximation of the complementary polynomial, supposing $Q_{0}$ is known. This is trivially the case if a $\delta>0$ is known with $||P(z)||_{\infty,\mathbb{T}}<1-\delta$; then $Q_{0}(z)=1$. In other cases, $Q_{0}(z)=1$ can be attained by first scaling $P(z)$ down slightly, $P(z)\to(1-\varepsilon)P(z)$ with a small $\varepsilon>0$.
First, for cases with a bound $||P(z)||_{\infty,\mathbb{T}}\leq(1-\delta)$, Algorithm 1 solves Problem 1 in time $O(N\log N)$ using the fast Fourier transform (FFT), where the even parameter $N\in\mathbb{Z}_{\geq d}$ defines the discrete Fourier basis $\{\mathrm{e}^{\mathrm{i}n\theta}\}_{n=-\frac{N}{2}+1}^{\frac{N}{2}}$. An informal description of our algorithm based on Corollary 2.2 with $Q_{0}(z)=1$ is as follows:
  1. 1.
    Compute approximations to the Fourier coefficients $(a_{n})_{n=-\frac{N}{2}+1}^{\frac{N}{2}}$ of $\log\big{(}1-\lvert P(\mathrm{e}^{\mathrm{i}\theta})\rvert^{2}\big{)}$ using the FFT, in time $O(N\log N)$.
  2. 2.
    Compute approximations to $Q$ at the $N$th roots of unity by applying the Fourier multiplier $\Pi$ in Fourier space, (1.6)–(1.8), using an FFT and inverse FFT in time $O(N\log N)$.
  3. 3.
    Compute approximations to the coefficients of $Q/{Q}_{0}$ in the monomial basis using the result of the previous step and the FFT, in time $O(N\log N)$.
We prove in Theorem 3 that this algorithm is efficient, with a sufficient $N$ asymptotically scaling as $N=O\!\left(\frac{d}{\delta}\log\frac{d}{\delta\varepsilon}\right)$.
In all other cases, when $||P(z)||_{\infty,\mathbb{T}}=1$ or $\delta$ is so small that the runtime $N\sim 1/\delta$ proven for Algorithm 1 is undesirable, Algorithm 2 solves Problem 1. It first scales down the input polynomial slightly before running Algorithm 1 in time $O(N\log N)$. Theorem 4 proves that a sufficient $N$ has asymptotic scaling $O\!\left(\frac{d}{\varepsilon}\log\frac{d}{\varepsilon}\right)$, eradicating any dependence on $\delta$.
We give a reference implementation of our algorithm and provide numerical evidence that it outperforms the optimization-based approach to complementary polynomials from [3] in Section 5.
1.2 Remarks on the results and their proofs
The following remarks apply to Theorem 2 and Corollary 2.2.
  1. 1.
    The crucial observation leading to Theorem 2 is that the real part of the function $\log(Q(z)/Q_{0}(z))$, where $Q$ is chosen so that all roots lie outside $\overline{\mathbb{D}}$ and the branch cuts are chosen appropriately, can be determined exactly on $\mathbb{T}$ using (1.1). This allows for the use of the Schwarz integral formula [17] to determine $Q$ on $\mathbb{D}$.
  2. 2.
    The representation (1.4c) of $Q$ is not manifestly a polynomial. Rather, as is elaborated in Section 2, it follows from (i) the existence of a solution of Problem 1 via the Féjer-Riesz theorem [18] and (ii) the uniqueness of analytic functions constructed by the Schwarz integral formula that $Q$ is a polynomial.
  3. 3.
    We show within the proof of Theorem 2 that, up to a multiplicative phase, the number of distinct solutions of Problem 1 is equal to $\prod_{j=1}^{d_{1}}(\beta_{j}+1)$, where $\beta_{j}$ is the multiplicity of the $j$th root ($j\in[d_{1}]$) of $Q$ outside of $\overline{\mathbb{D}}$. However, constructing all of these solutions requires knowledge of all roots of $1-P(z)P^{*}(1/z)$ on $\mathbb{C}$. We construct a canonical solution of Problem 1 with no roots on $\mathbb{D}$ in (1.4c); this requires only the knowledge of the roots of $1-\lvert P(z)\rvert^{2}$ on $\mathbb{T}$.
  4. 4.
    Theorem 2 and Corollary 2.2 can be proven in different ways. Consider the function $\log(Q(z)/Q_{0}(z))$, where $Q$ is chosen so that all roots lie outside $\overline{\mathbb{D}}$ and the branch cuts are chosen appropriately. One can use (1.1) and the Fejér-Riesz theorem to construct a scalar Riemann-Hilbert problem [19] on $\mathbb{T}$ for $\log(Q(z)/Q_{0}(z))$; this Riemann-Hilbert problem is explicitly solvable using a Cauchy integral, from which (1.4a) follows. Corollary 2.2 can be proven directly using (1.1), the Fejér-Riesz theorem, and the fact that the periodic Hilbert transform (2.11) relates the real and imaginary parts of the boundary values of a function analytic on $\mathbb{D}$, namely $\log(Q(z)/Q_{0}(z))$.
1.3 Related work
In the quantum signal processing literature, a variety of numerical methods for solving Problem 1 or its avatars have been developed. As is evident from the proof of Theorem 2 in Section 2.1, knowledge of all roots of $1-P(z)P^{*}(1/z)$ allows for the explicit construction of $Q$; analogous statements hold for complementary polynomials in standard QSP. Thus, employing standard root-finding algorithms provides a straightforward means to calculate complementary polynomials [4, 11, 13]. Unfortunately, root-finding algorithms are known to be expensive and suffer from numerical instability, and the highest degree polynomial successfully treated with this approach was reported to have $d=3\times 10^{3}$ [13]. An alternative method that avoids root-finding and instead directly calculates the characteristic polynomial of the roots of $1-P(z)P^{*}(1/z)$ within $\mathbb{D}$ using Prony’s method has been proposed in [15]. Numerical experiments have demonstrated the effectiveness of this approach for polynomials with degree up to $d=5\times 10^{4}$.
The current state-of-the-art method for Problem 1 was developed in [3]. There, a loss function derived from the complementarity condition (1.1) is minimized to determine $Q$ with an optimization procedure; this approach was demonstrated to be effective for $d$ up to the order of $10^{7}$, achieving accuracies as low as $10^{-6}$ in the loss function. In this paper, we present numerical results showing that our algorithm is effective for the same degrees, up to $d=10^{7}$. At the same time our algorithm requires much shorter runtimes and achieves better accuracies, even without the GPU acceleration used in [3].
We mention in passing that as an alternative to the approaches described above, optimization-based methods to compute phase factors without knowledge of the complementary polynomial have been developed in [14, 20, 21]. These methods have been used to determine phase factors for polynomials up to degree $d=10^{5}$.
1.4 Notation
We write the complex conjugate of $z\in\mathbb{C}$ as $z^{*}$ and for a function $f:\mathbb{C}\to\mathbb{C}$, define $f^{*}(z)\coloneqq f(z^{*})^{*}$. For a set $X\subset\mathbb{C}$, we write $\partial X$ and $\overline{X}$ for its boundary and closure, respectively. We define $\mathbb{D}$ and $\mathbb{T}$ to be open unit disk and unit circle in the complex plane, respectively. Given an integer $N\in\mathbb{Z}_{>0}$, we define the sets $[N]\coloneqq\{n\in\mathbb{Z}_{>0}:n\leq N\}$ and $[N]_{0}\coloneqq[N]\cup\{0\}$. Dashed integrals indicate a principal value prescription with respect to singularities of the integrand on the integration contour. Contour integrals are always assumed to carry a positive orientation. Unless otherwise indicated, all logarithms are with respect to base $\mathrm{e}$. We denote by $\lVert\cdot\rVert_{\infty,\mathbb{T}}$ the uniform norm on $\mathbb{T}$.
1.5 Plan of the paper
Theorem 2 and Corollaries 2.1 and 2.2 are proved in Section 2. In Section 3, we design two algorithms for the computation of $Q$ following from Corollary 2.2. Error analysis of the algorithms is performed in Section 4 and numerical results comparing our algorithm to the optimization-based approach of [3] are presented in Section 5. Section 6 contains a discussion of our results and possibilities for future work. In Appendix A, we show that different QSP parameterizations can be viewed as special cases of GQSP.
2 Proofs of main results
We provide rigorous proofs of the mathematical results reported in the previous section. Theorem 2 is proved in Section 2.1 and Corollaries 2.1 and 2.2 are proved in Sections 2.2 and 2.3, respectively.
2.1 Proof of Theorem 2
Observe that on $\mathbb{T}$, $1-\lvert P(z)\rvert^{2}=1-P(z)P^{*}(1/z)$, a positive-semidefinite Laurent polynomial of degree $d$. Thus, by the Fejér-Riesz theorem, there exists $Q\in\mathbb{C}[z]$ so that $\deg Q=d$, $Q$ is nonzero on $\mathbb{D}$, any root of $Q$ on $\mathbb{T}$ has even multiplicity, and (1.1) is satisfied. We write
$Q(z)=\bar{Q}\Bigg{(}\prod_{j=1}^{d_{0}}(z-t_{j})^{\alpha_{j}}\Bigg{)}\Bigg{(}\prod_{j=1}^{d_{1}}(z-w_{j})^{\beta_{j}}\Bigg{)}\quad(z\in\mathbb{C}),$ (2.1)
where $\bar{Q}\in\mathbb{C}\setminus\{0\}$ and $\{(w_{j},\beta_{j})\}_{j=1}^{d_{1}}$ are the roots of $Q$ outside of $\mathbb{D}$ with corresponding multiplicities; recall that $2\alpha_{j}$ is the multiplicity of the root $t_{j}$. It follows from (1.1) and (2.1) that
$1-\lvert P(z)\rvert^{2}=\lvert Q(z)\rvert^{2}=\lvert\bar{Q}\rvert^{2}\Bigg{(}\prod_{j=1}^{d_{0}}(z-t_{j})^{\alpha_{j}}\bigg{(}\frac{1}{z}-\frac{1}{t_{j}}\bigg{)}^{\alpha_{j}}\Bigg{)}\Bigg{(}\prod_{j=1}^{d_{1}}(z-w_{j})^{\beta_{j}}\bigg{(}\frac{1}{z}-w_{j}^{*}\bigg{)}^{\beta_{j}}\Bigg{)}\quad(z\in\mathbb{T});$ (2.2)
note that this factorization is unique up to rotations $\bar{Q}\to t\bar{Q}$, $t\in\mathbb{T}\simeq\mathrm{U}(1)$. Moreover, we see that transforming
$Q(z)\to\bigg{(}\frac{1-zw_{j}^{*}}{z-w_{j}}\bigg{)}^{k}Q(z)$ (2.3)
for any $j\in[d_{1}]$ and $k\in[\beta_{j}]_{0}$, preserves (2.2) via (1.1). It follows that there are $\prod_{j=1}^{d_{1}}(\beta_{j}+1)$ distinct solutions of Problem 1, up to $\mathrm{U}(1)$ equivalence.
To construct a canonical solution (2.1) of Problem 1, we combine (1.5) and (2.1) and write
$\frac{Q(z)}{{Q}_{0}(z)}=\bar{Q}\prod_{j=1}^{d_{1}}(z-w_{j})^{\beta_{j}}.$ (2.4)
Observe that any logarithm of $Q/Q_{0}$ will have branch points $\{w_{j}\}_{j=1}^{d_{1}}$. Consider the function
$U(z)\coloneqq\log\bigg{(}\frac{Q(z)}{{Q}_{0}(z)}\bigg{)}\quad(z\in\mathbb{C}\setminus B),$ (2.5)
where the branch cuts are chosen to be
$B=\bigcup_{n=1}^{d_{1}}\{sw_{n}:s\in[1,\infty)\}.$ (2.6)
By construction, $U(z)$ is analytic on $\overline{\mathbb{D}}$. The real part of $U(z)$ is found to be
$\mathrm{Re}\,U(z)=\log\bigg{\lvert}\frac{Q(z)}{Q_{0}(z)}\bigg{\rvert}=\frac{1}{2}\log\bigg{(}\frac{1-\lvert P(z)\rvert^{2}}{\lvert Q_{0}(z)\rvert^{2}}\bigg{)}\quad(z\in\mathbb{C}\setminus B).$ (2.7)
In particular, (2.7) holds on $\mathbb{T}$, so by the Schwarz integral formula [17], we obtain a representation of $U(z)$ on $\mathbb{D}$,
$U(z)=$ $\;\frac{1}{2\pi\mathrm{i}}\int_{\mathbb{T}}\frac{z^{\prime}+z}{z^{\prime}-z}\,\mathrm{Re}\,U(z^{\prime})\,\frac{\mathrm{d}z^{\prime}}{z^{\prime}}+\mathrm{i}\,\mathrm{Im}\,U(0)$
$=$ $\;\frac{1}{4\pi\mathrm{i}}\int_{\mathbb{T}}\frac{z^{\prime}+z}{z^{\prime}-z}\log\bigg{(}\frac{1-\lvert P(z^{\prime})\rvert^{2}}{\lvert Q_{0}(z^{\prime})\rvert^{2}}\bigg{)}\,\frac{\mathrm{d}z^{\prime}}{z^{\prime}}+\mathrm{i}\,\mathrm{Im}\,U(0)\quad(z\in\mathbb{D}).$ (2.8)
By exponentiating (2.5) and (2.1) and using the $\mathrm{U}(1)$ symmetry of Problem 1, we obtain (1.4a).
The second case of (1.4c) is obtained from the first using the Plemelj formula [19]. Note that the integrand in (1.4a) has a simple pole at $z^{\prime}=z$ with residue $2\log\big{(}\frac{1-\lvert P(z)\rvert^{2}}{\lvert Q_{0}(z)\rvert^{2}}\big{)}$. Thus, applying the Plemelj formula as $z\in\mathbb{D}$ approaches the contour $\mathbb{T}$ gives (1.4b).
The third case of (1.4c) is obtained from the second by analytic continuation. Let us write (1.4b) as
$Q(z)=$ $\;Q_{0}(z)\exp\Bigg{(}\frac{1}{4\pi\mathrm{i}}\,\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{T}}\frac{z^{\prime}+z}{z^{\prime}-z}\log\bigg{(}\frac{1-\lvert P(z^{\prime})\rvert^{2}}{\lvert{Q}_{0}(z^{\prime})\rvert^{2}}\bigg{)}\frac{\mathrm{d}z^{\prime}}{z^{\prime}}-\frac{1}{2}\log\bigg{(}\frac{1-\lvert P(z)\rvert^{2}}{\lvert{Q}_{0}(z)\rvert^{2}}\bigg{)}+\log\bigg{(}\frac{1-\lvert P(z)\rvert^{2}}{\lvert{Q}_{0}(z)\rvert^{2}}\bigg{)}\Bigg{)}$
$=$ $\;Q_{0}(z)\frac{1-\lvert P(z)\rvert^{2}}{\lvert Q_{0}(z)\rvert^{2}}\exp\Bigg{(}\frac{1}{4\pi\mathrm{i}}\,\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{T}}\frac{z^{\prime}+z}{z^{\prime}-z}\log\bigg{(}\frac{1-\lvert P(z^{\prime})\rvert^{2}}{\lvert{Q}_{0}(z^{\prime})\rvert^{2}}\bigg{)}\frac{\mathrm{d}z^{\prime}}{z^{\prime}}-\frac{1}{2}\log\bigg{(}\frac{1-\lvert P(z)\rvert^{2}}{\lvert{Q}_{0}(z)\rvert^{2}}\bigg{)}\Bigg{)}\quad(z\in\mathbb{T}).$ (2.9)
Analytic continuation of the prefactor and exponent in (2.1), using that the latter represents the boundary values of a Cauchy integral, to $\mathbb{C}\setminus\overline{\mathbb{D}}$ gives (1.4c).
2.2 Proof of Corollary 2.1
This result follows from the Féjer-Riesz theorem and properties of Laurent polynomials with real coefficients.
On $\mathbb{T}$, $1-\lvert P(z)\rvert^{2}=1-P(z)P(1/z)$, a positive-semidefinite Laurent polynomial of degree $d$ with real coefficients. By the same argument as in the proof of Theorem 2, we may write $Q$ in the canonical form (2.1). Because $1-P(z)P(1/z)$ has real coefficients, it has the following symmetries (i) if $w\in\mathbb{C}\setminus(\mathbb{R}\cup\mathbb{T})$ is a root, so are $1/w$, $w^{*}$, and $1/w^{*}$ and (ii) if $w\in\mathbb{T}\setminus\{\pm 1\}$ is a root, so is $1/w$, in both cases with the same multiplicities. Requiring that these symmetries be respected in (2.2) shows that the polynomial $Q(z)/\bar{Q}$ obtained from (2.1) has real coefficients. Choosing $\bar{Q}\in\mathbb{R}$ gives the result.
2.3 Proof of Corollary 2.2
Performing the change of variables $z=\mathrm{e}^{\mathrm{i}\theta}$, $z^{\prime}=\mathrm{e}^{\mathrm{i}\theta^{\prime}}$ in (1.4b) gives
$Q(\mathrm{e}^{\mathrm{i}\theta})=Q_{0}(\mathrm{e}^{\mathrm{i}\theta})\exp\Bigg{(}\frac{1}{4\pi\mathrm{i}}\,\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{-\pi}^{\pi}\cot\bigg{(}\frac{\theta^{\prime}-\theta}{2}\bigg{)}\log\bigg{(}\frac{1-\lvert P(\mathrm{e}^{\mathrm{i}\theta^{\prime}})\rvert^{2}}{\lvert Q_{0}(\mathrm{e}^{\mathrm{i}\theta^{\prime}})\rvert^{2}}\bigg{)}\,\mathrm{d}\theta^{\prime}+\frac{1}{2}\log\bigg{(}\frac{1-\lvert P(\mathrm{e}^{\mathrm{i}\theta})\rvert^{2}}{\lvert Q_{0}(\mathrm{e}^{\mathrm{i}\theta})\rvert^{2}}\bigg{)}\Bigg{)}.$ (2.10)
The first term in the exponent is identified as a periodic Hilbert transform [22],
$H[f(\theta)]\coloneqq\frac{1}{2\pi}\,\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{-\pi}^{\pi}\cot\bigg{(}\frac{\theta^{\prime}-\theta}{2}\bigg{)}f(\theta^{\prime})\,\mathrm{d}\theta^{\prime}.$ (2.11)
We recall that the periodic Hilbert transform (2.11) has the complex exponentials as eigenfunctions,
$H[\mathrm{e}^{\mathrm{i}n\theta}]=\begin{cases}\mathrm{i}\mathrm{e}^{\mathrm{i}n\theta}&n\in\mathbb{Z}_{>0}\\ 0&n=0\\ -\mathrm{i}\mathrm{e}^{\mathrm{i}n\theta}&n\in\mathbb{Z}_{<0}.\end{cases}$ (2.12)
Writing $\Pi=\frac{1}{2}(1-\mathrm{i}H)$, we see from (2.12) that (1.7) holds. Expressing (2.10) in terms of $\Pi$ gives the result.
3 Numerical methods
We develop a numerical method for solving Problem 1 based on Corollary 2.2 in the case $Q_{0}(z)=1$. Our starting point is the Laurent series
$S(z)\coloneqq\sum_{n\in\mathbb{Z}}a_{n}z^{n},$ (3.1)
where
$a_{n}\coloneqq\frac{1}{2\pi\mathrm{i}}\int_{\mathbb{T}}\log\big{(}{1-\lvert P(z)\rvert^{2}}\big{)}\frac{\mathrm{d}z}{z^{n+1}}\quad(n\in\mathbb{Z}).$ (3.2)
Observe that $S(\mathrm{e}^{\mathrm{i}\theta})$ is the Fourier series of $\log\big{(}{1-\lvert P(\mathrm{e}^{\mathrm{i}\theta})\rvert^{2}}\big{)}$. Thus, from (1.6) and (1.7), we have
$Q(\mathrm{e}^{\mathrm{i}\theta})=\exp\big{(}\Pi[{S}(\mathrm{e}^{\mathrm{i}\theta})]\big{)}=\exp\Bigg{(}\frac{1}{2}a_{0}+\sum_{n=1}^{\infty}a_{n}\mathrm{e}^{\mathrm{i}n\theta}\Bigg{)}.$ (3.3)
To numerically evaluate (3.3), we make two approximations that allow us to compute $Q$ in the monomial basis by a sequence of FFTs. As the performance of an FFT is optimized when the number of Fourier modes is a power of $2$, we choose the size of this basis to be $N=2^{M}$ for some $M\in\mathbb{Z}_{\geq 1}$ satisfying $M\geq\lceil\log_{2}(d+1)\rceil$. The error analysis of Algorithm 1, which will result from the approximations and analysis in this section, is performed in Section 4. First, we introduce the Laurent polynomial truncation of (3.1)
$S_{N}(z)\coloneqq\sum_{n=-\frac{N}{2}+1}^{\frac{N}{2}}a_{n}z^{n}\quad(N\in\mathbb{Z}_{\geq d_{1}+1}).$ (3.4)
Second, we will approximate the coefficients (3.2) by discrete Fourier transforms. Consider the root of unity
$\omega_{N}\coloneqq\mathrm{e}^{2\pi\mathrm{i}/N}.$ (3.5)
which we use to define the following approximation of the Laurent coefficients (3.2),
$\tilde{a}_{n}\coloneqq\frac{1}{N}\sum_{m=-\frac{N}{2}+1}^{\frac{N}{2}}\log\big{(}{1-\lvert P(\omega_{N}^{m})\rvert^{2}}\big{)}\omega_{N}^{-nm}\quad(n=-\tfrac{N}{2}+1,\ldots,\tfrac{N}{2}).$ (3.6)
It follows that
$\tilde{S}_{N}(z)\coloneqq\sum_{n=-\frac{N}{2}+1}^{\frac{N}{2}}\tilde{a}_{n}z^{n}$ (3.7)
is an approximation of (3.4).
Replacing $S(\mathrm{e}^{\mathrm{i}\theta})$ by $\tilde{S}_{N}(\mathrm{e}^{\mathrm{i}\theta})$ in (3.3) gives
$\tilde{Q}_{1,N}(\mathrm{e}^{\mathrm{i}\theta})\coloneqq\exp\big{(}\Pi\big{[}\tilde{S}_{N}(\mathrm{e}^{\mathrm{i}\theta})\big{]}\big{)}=\exp\Bigg{(}\frac{1}{2}\tilde{a}_{0}+\sum_{n=1}^{\frac{N}{2}}\tilde{a}_{n}\mathrm{e}^{\mathrm{i}n\theta}\Bigg{)}$ (3.8)
as an approximation of $Q$ on $\mathbb{T}$. This is, however, no guarantee that $\tilde{Q}_{1,N}(\mathrm{e}^{\mathrm{i}\theta})$ is a trigonometric polynomial or equivalently, extends to a polynomial $\tilde{Q}_{1,N}(z)$ on $\mathbb{C}$. We can instead (i) interpolate (3.8) through the roots of unity $\{\omega_{N}^{n}\}_{n=0}^{N-1}$ and (ii) discard terms in $z^{n}$ for $n>d$ to obtain an explicit polynomial of degree $d$ in the monomial basis.
To this end, let us write
$Q(z)=\sum_{n=0}^{d}q_{n}z^{n},$ (3.9)
in correspondence with (1.3). We define $q_{n}=0$ for $n>d$. The evaluation of $Q$ at the roots of unity $\{\omega_{N}^{n}\}_{n=0}^{N-1}$,
$Q(\omega_{N}^{n})=\sum_{m=0}^{N-1}q_{m}\omega_{N}^{nm}\quad(n\in[N-1]_{0}),$ (3.10)
is an inverse discrete Fourier transform of the coefficients $(q_{n})_{n=0}^{N-1}$. Thus, the corresponding forward transform allows for the computation of $(q_{n})_{n=0}^{d}$,
$q_{n}=\frac{1}{N}\sum_{m=0}^{N-1}Q(\omega_{N}^{m})\omega_{N}^{-nm}\quad(n\in[d]_{0}).$ (3.11)
We are led to define the following approximations of the monomial coefficients $(q_{n})_{n=0}^{d}$,
$\tilde{q}_{n}\coloneqq\frac{1}{N}\sum_{m=0}^{N-1}\tilde{Q}_{1,N}(\omega_{N}^{m})\omega_{N}^{-nm}\quad(n\in[d]_{0})$ (3.12)
and the following manifestly polynomial approximation to $Q$,
$\tilde{Q}_{2,N}(z)\coloneqq\sum_{n=0}^{d}\tilde{q}_{n}z^{n}.$ (3.13)
3.1 Algorithm
We combine the observations obtained in this section into the following algorithms to compute $\tilde{Q}_{2,N}(z)$ (3.13), the numerical approximation to the canonical complementary polynomial to $P(z)$. In order to avoid costly root-finding to determine $Q_{0}(z)$, we will consider situations where $Q_{0}(z)=1$, see Algorithm 1. However, this is not a restriction; by scaling down the input polynomial ever so slightly, $Q_{0}(z)=1$ can always be achieved, leading to the general Algorithm 2 applicable to any polynomial.
For many practical applications of QSP-type algorithms in quantum computation, a parameter $\delta>0$ with
$\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 1-\delta$ (3.14)
can be controlled a priori in the construction of the polynomial $P(z)$ approximating a target function. Then, (3.14) ensures that $1-|P(z)|^{2}$ has no roots on the unit circle $z\in\mathbb{T}$, and $Q_{0}(z)=1$. The numerical method explained above can then be used directly, and is summarised in the following Algorithm 1:
Algorithm 1 (Construction of the complementary polynomial for known $\delta$)
Input:
  • The coefficients (in the monomial basis) of a polynomial $P\in\mathbb{C}[z]$ of degree $d$ with $\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 1-\delta$, for a $\delta>0$.
  • An integer $N\geq d$ controlling the accuracy $\varepsilon$ of the output, called dimension of the FFT.
Output:
  • The coefficients (in the monomial basis) of the canonical complementary polynomial $Q\in\mathbb{C}[z]$ of degree $d$ from Theorem 2, up to accuracy $\varepsilon$.
Complexity:
  • Runtime: $O(N\log N)$
  • Sufficient $N$ for accuracy $\varepsilon$: $N=O\!\left(\frac{d}{\delta}\log\frac{d}{\delta\varepsilon}\right)$, see Theorem 3
Algorithm:
  1. 1.
    Compute $(P(\omega_{N}^{n}))_{n=0}^{N-1}$, the input polynomial $P(z)=\sum_{n=0}^{d}p_{n}z^{n}$ evaluated at all $N$th roots of unity, with the inverse FFT of $(p_{n})_{n=0}^{N-1}$ padded with zeros, i.e. $p_{n}=0,n>d$.
  2. 2.
    Compute $(\tilde{a}_{n})_{n=-\frac{N}{2}+1}^{\frac{N}{2}}$ by applying the FFT to $(\log\left(1-\left|P(\omega_{N}^{n})\right)|^{2}\right))_{n=0}^{N-1}$ obtained from the previous step, see (3.6).
  3. 3.
    Apply the Fourier multiplier $\Pi$ to the truncated Fourier series (3.7).
  4. 4.
    Evaluate the exponential $(\tilde{Q}_{1,N}(\omega_{N}^{n}))_{n=0}^{N-1}$ in (3.8) at $N$th roots of unity by taking the exponential of an inverse FFT of the previous step’s result.
  5. 5.
    Compute $(\tilde{q}_{n})_{n=0}^{N-1}$ in (3.12) by applying the FFT to $(\tilde{Q}_{1,N}(\omega_{N}^{n}))_{n=0}^{N-1}$.
  6. 6.
    Truncate the coefficients of the previous step to $(\tilde{q}_{n})_{n=0}^{d}$ and output them as the coefficients of the approximation $\tilde{Q}_{2,N}(z)$ to the complementary polynomial, see (3.13).
Reference implementation:
  • See Figure 1 for Python code, and Section 3 for numerical results.
The algorithm relies heavily on the fast Fourier transform (FFT) to switch between coefficients of polynomials and values of polynomials at roots of unity, and to apply the Fourier multiplier in Fourier space. The FFTs of dimension $N$ result in the runtime $O(N\log N)$.
In Theorem 3 in the next section, we prove that Algorithm 1 computes the canonical complementary polynomial (the unique solution, up to phase, with no roots in $\mathbb{D}$, as in Theorem 2), up to accuracy $\varepsilon$ in the coefficients. Moreover, the same theorem shows that the algorithm is efficient in degree and error because of its asymptotic scaling $N=O(d\log(d/\varepsilon))$.
When $\delta$ is very small, the scaling $1/\delta$ from the bound on $N$ in Theorem 3 suggests a long runtime; in the extreme case $\delta=0$, the proof for Algorithm 1 fails because $Q_{0}(z)\neq 1$. For those cases, we present Algorithm 2, in which the initial polynomial is first scaled down as $P(z)\to(1-\tfrac{\varepsilon}{4})P(z)$ to achieve an effective $\delta=\varepsilon/4$. We will prove in Theorem 4 that the polynomial generated by Algorithm 2 with asymptotic scaling $N=O(d/\varepsilon\,\log(d/\varepsilon))$ satisfies the complementarity condition up to accuracy $\varepsilon$, i.e.,
$\big{\lVert}\lvert P(z)\rvert^{2}+\lvert Q(z)\rvert^{2}-1\big{\rVert}_{\infty,\mathbb{T}}\leq\varepsilon.$ (3.15)
Closeness in complementarity condition is a weaker statement than closeness to the exact canonical complementary polynomial like the output of Algorithm 1. Yet, it enables us to rigorously and efficiently extend our numerical method to all polynomials:
Algorithm 2 (Construction of the complementary polynomial for unknown, small, or zero $\delta$)
Input:
  • The coefficients (in the monomial basis) of a polynomial $P\in\mathbb{C}[z]$ of degree $d$ with $|P(z)|\leq 1$, $z\in\mathbb{T}$.
  • An integer $N\geq d$ controlling the accuracy $\varepsilon$ of the output, called dimension of the FFT.
Output:
  • The coefficients (in the monomial basis) of a polynomial $Q\in\mathbb{C}[z]$ of degree $d$ that satisfies the complementarity condition up to accuracy $\varepsilon$, i.e., $\big{\lVert}|P(z)|^{2}+|Q(z)|^{2}-1\big{\rVert}_{\infty,\mathbb{T}}\leq\varepsilon$.
Complexity:
  • Runtime: $O(N\log N)$
  • Sufficient $N$ for accuracy $\varepsilon$: $N=O\!\left(\frac{d}{\varepsilon}\log\frac{d}{\varepsilon}\right)$, see Theorem 4
Algorithm:
  1. 1.
    Compute $\big{(}(1-\tfrac{\varepsilon}{4})p_{n}\big{)}_{n=0}^{d}$ to scale down the input polynomial $P(z)=\sum_{n=0}^{d}p_{n}z^{n}$.
  2. 2.
    Return the result of Algorithm 1 with input the downscaled polynomial from the previous step, for which $\delta=\varepsilon/4$, and $N$ chosen to yield an accuracy of $\varepsilon/6(d+1)$.
Remark 3.1.1.
Empirically, we find that Algorithm 2 may not be needed. Even without the initial downscaling, our numerical results in Section 5 suggest that Algorithm 1 alone is efficacious even when $\delta=0$, with error dependence $N\approx O(d/\sqrt[4]{\varepsilon})$. Here, and in our error analysis in the next section, we prefer to include Algorithm 2 to preserve full mathematical rigor.
4 Error analysis of algorithms
We perform error analysis on Algorithm 1 and Algorithm 2, developed in the previous section. We state our main results, Theorem 3, Corollary 3.1, and Theorem 4 below. The corresponding proofs are given in Sections 4.3, 4.4, and 4.5, respectively.
4.1 Error metrics
We introduce two error metrics that we will later use to analyze the algorithms. In the definitions of these error metrics, we view $P,Q\in\mathbb{C}[z]$ as a priori unrelated; when $Q$ is an exact complementary polynomial to a given $P$, the error metrics will evaluate to zero.
The first error metric we consider is motivated by the complementarity condition (1.1),
$\Phi(P,Q)\coloneqq\big{\lVert}\lvert P(z)\rvert^{2}+\lvert Q(z)\rvert^{2}-1\big{\rVert}_{\infty,\mathbb{T}}.$ (4.1)
The second error metric we consider was introduced in [3] as a loss function for optimization of $Q$ and is defined in terms of the monomial coefficients of the polynomials $P$ and $Q$,
$\tilde{\Phi}(P,Q)\coloneqq\Bigg{(}\sum_{n=-d}^{d}\Bigg{\lvert}\sum_{m=0}^{d}\big{(}p_{n+m}p_{m}^{*}+{q}_{n+m}{q}_{m}^{*}\big{)}-{\delta_{n,0}}\Bigg{\rvert}^{2}\Bigg{)}^{\frac{1}{2}}.$ (4.2)
The error metrics (4.1) and (4.2) are compatible in the following sense.
Proposition 4.1.1.
Let $P,Q\in\mathbb{C}[z]$ with $\deg P=\deg Q=d$. Then, the complementarity condition $\Phi(P,Q)$ and the loss function $\tilde{\Phi}(P,Q)$ are equivalent in the sense that they satisfy the inequalities
$\frac{1}{\sqrt{2d+1}}\Phi(P,Q)\leq\tilde{\Phi}(P,Q)\leq\sqrt{2d+1}\Phi(P,Q).$ (4.3)
Proof.
Putting (1.3) and (3.9) into (4.1), we write
$\Phi(P,Q)=\Bigg{\lVert}\sum_{n,m=1}^{d}\big{(}p_{n}p_{m}^{*}+q_{n}q_{m}^{*}\big{)}z^{n-m}-1\Bigg{\rVert}_{\infty,\mathbb{T}}.$ (4.4)
By changing the summation variables and using the triangle and $\ell^{1}$-$\ell^{2}$ norm inequalities, we obtain
$\Phi(P,Q)=$ $\;\Bigg{\lVert}\sum_{n=-d}^{d}\Bigg{(}\sum_{m=0}^{d}(p_{n+m}p_{m}^{*}+q_{n+m}q_{m}^{*}\big{)}-\delta_{n,0}\Bigg{)}z^{n}\Bigg{\rVert}\leq\sum_{n=-d}^{d}\Bigg{\lvert}\sum_{m=0}^{d}(p_{n+m}p_{m}^{*}+q_{n+m}q_{m}^{*}\big{)}-\delta_{n,0}\Bigg{\rvert}$
$\leq$ $\;\sqrt{2d+1}\Bigg{(}\sum_{n=-d}^{d}\Bigg{\lvert}\sum_{m=0}^{d}(p_{n+m}p_{m}^{*}+q_{n+m}q_{m}^{*}\big{)}-\delta_{n,0}\Bigg{\rvert}^{2}\Bigg{)}^{\frac{1}{2}},$ (4.5)
which, recalling (4.2), is the first inequality in (4.3).
To prove the second inequality in (4.3), we use Cauchy integral formula to write
$\sum_{m=0}^{d}\big{(}p_{n+m}p_{m}^{*}+q_{n+m}q_{m}^{*}\big{)}-\delta_{n,0}=\frac{1}{2\pi\mathrm{i}}\int_{\mathbb{T}}\Bigg{(}\sum_{l=-d}^{d}\sum_{m=0}^{d}(p_{l+m}p_{m}^{*}+q_{l+m}q_{m}^{*}-\delta_{l,0})z^{n}\Bigg{)}\frac{\mathrm{d}z}{z^{n+1}}\quad(n=-d,\ldots,d)$ (4.6)
and hence,
$\Bigg{\lvert}\sum_{m=0}^{d}\big{(}p_{n+m}p_{m}^{*}+q_{n+m}q_{m}^{*}\big{)}-\delta_{n,0}\Bigg{\rvert}\leq\Phi(P,Q)\quad(n=-d,\ldots,d),$ (4.7)
where we have used (4.1). Putting (4.7) into (4.2) gives the result. ∎
4.2 Results of error analysis
Here, we establish rigorous error bounds on Algorithms 1 and 2, using the error metrics introduced in the previous subection.
Theorem 3 (Error bounds for Algorithm 1).
Suppose $P\in\mathbb{C}[z]$ satisfying $\lVert P(z)\rVert_{\mathbb{T},\infty}\leq 1-\delta$ for some $\delta\in(0,1)$ and $\varepsilon\in(0,1)$ are given. Choose $N\in\mathbb{Z}_{\geq 1}$ such that
$N\geq N_{0}(\varepsilon,\delta,d)\coloneqq\bigg{\lceil}\frac{2}{\log r_{\delta}}\log\bigg{(}8\frac{\log(\frac{1}{\delta})}{r_{\delta}-1}\frac{1}{\varepsilon}\bigg{)}\bigg{\rceil},$ (4.8)
where
$r_{\delta}\coloneqq\bigg{(}\frac{1}{1-\delta}\bigg{)}^{\frac{1}{d}}.$ (4.9)
Then, the output of Algorithm 1 satisfies
$\lvert q_{n}-\tilde{q}_{n}\rvert<\varepsilon\quad(n\in[d]_{0}).$ (4.10)
In particular, (4.8) has the joint asymptotic complexity
$N_{0}(\varepsilon,\delta,d)=O\bigg{(}\frac{d}{\delta}\log\frac{d}{\delta\varepsilon}\bigg{)}.$ (4.11)
We can use Theorem 3 to bound the error metrics introduced in Section 4.1
Corollary 3.1.
Suppose that (4.10) holds for some $\varepsilon\in(0,1)$. Then, the error metrics (4.1) and (4.2) satisfy the inequalities
$\Phi(P,\tilde{Q}_{2,N})<(d+1)(d+3)\varepsilon$ (4.12)
and
$\tilde{\Phi}(P,\tilde{Q}_{2,N})<3(d+1)(2d+1)\varepsilon.$ (4.13)
Letting $\varepsilon^{\prime}=\tilde{\Phi}(P,\tilde{Q}_{2,N})$, we see from Theorem 3 that for fixed $\delta$, $N=O(d\log(1/\varepsilon^{\prime}))$ is required to achieve (4.13). Numerical results verifying this assertion are presented in Section 5.
Theorem 4 (Error bounds for Algorithm 2).
Suppose $P\in\mathbb{C}[z]$ satisfying $\lVert P(z)\rVert_{\mathbb{T},\infty}\leq 1$ and $\varepsilon\in(0,1)$ are given. Choose $N\in\mathbb{Z}_{\geq 1}$ such that
$N\geq N_{0}\bigg{(}\frac{\varepsilon}{4},\frac{\varepsilon}{5(d+1)},d\bigg{)}$ (4.14)
with $N_{0}$ defined in (4.8). Then, the result $\tilde{Q}_{2,N}\in\mathbb{C}[z]$ of Algorithm 2 satisfies the bound
$\Phi(P,\tilde{Q}_{2,N})<\varepsilon$ (4.15)
on the complementarity condition (4.1). In particular, we have the joint asymptotic complexity
$N_{0}\bigg{(}\frac{\varepsilon}{4},\frac{\varepsilon}{5(d+1)},d\bigg{)}=O\bigg{(}\frac{d}{\varepsilon}\log\frac{d}{\varepsilon}\bigg{)}.$ (4.16)
4.3 Proof of Theorem 3
Let
$R\coloneqq\min_{j\in[d_{1}]}\lvert w_{j}\rvert$ (4.17)
and define the function
$M(r)\coloneqq\max_{\rho=1/r,r}\max_{z\in\mathbb{T}}\big{\lvert}\log\big{(}{1-P(\rho z)P^{*}(1/\rho z)}\big{)}\big{\rvert}\quad(r\in(1,R)).$ (4.18)
Our analysis is based on the following lemma.
Lemma 4.3.1.
For $r\in(1,R)$, the Fourier coefficients $(a_{n})_{n\in\mathbb{Z}}$ from (3.2) satisfy
$\lvert a_{n}\rvert\leq M(r)r^{-\lvert n\rvert}\quad(n\in\mathbb{Z}).$ (4.19)
Proof.
For any $r\in(1,R)$, the function $\log\big{(}{1-\lvert P(z)\rvert^{2}}\big{)}$ may be analytically continued to the closure of the annulus
$A(r)\coloneqq\{z\in\mathbb{C}:\tfrac{1}{r}<\lvert z\rvert<r\}.$ (4.20)
Suppose $n\in\mathbb{Z}_{\geq 0}$. Then, using Cauchy’s theorem to deform the contour in (3.2), we find
$\lvert a_{n}\rvert=$ $\;\frac{1}{2\pi}\Bigg{\lvert}\int_{\lvert z\rvert=r}\log\big{(}{1-P(z)P^{*}(1/z)}\big{)}\frac{\mathrm{d}z}{z^{n+1}}\Bigg{\rvert}=r^{-n}\frac{1}{2\pi}\Bigg{\lvert}\int_{\mathbb{T}}\log\big{(}{1-P(rz)P^{*}(1/rz)}\big{)}\frac{\mathrm{d}z}{z^{n+1}}\Bigg{\rvert}$
$\leq$ $\;r^{-n}\frac{1}{2\pi}\int_{\mathbb{T}}\bigg{\lvert}\log\big{(}{1-P(rz)P^{*}(1/rz)}\big{)}\frac{1}{z^{n+1}}\Bigg{\rvert}\,\mathrm{d}z=r^{-n}\frac{1}{2\pi}\int_{\mathbb{T}}\big{\lvert}\log\big{(}{1-P(rz)P^{*}(1/rz)}\big{)}\big{\rvert}\,\mathrm{d}z\leq L(r)r^{-n},$ (4.21)
where
$L(r)\coloneqq\sup_{\frac{1}{r}<\rho<r}\frac{1}{2\pi}\int_{\mathbb{T}}\big{\lvert}\log\big{(}{1-P(\rho z)P^{*}(1/\rho z)}\big{)}\big{\rvert}\,\mathrm{d}z\quad(r\in(1,R)),$ (4.22)
for each $n\in\mathbb{Z}_{\geq 0}$. A similar argument for $n\in\mathbb{Z}_{\leq 0}$ shows that
$\lvert a_{n}\rvert\leq L(r)r^{n}\quad(n\in\mathbb{Z}_{\leq 0}).$ (4.23)
By the maximum modulus principle, we have
$L(r)\leq M(r)\quad(r\in(1,R)).$ (4.24)
The result (4.19) follows by combining (4.3) and (4.23) with (4.24). ∎
Using Lemma 4.3.1, we readily obtain a bound on the truncation error,
$\big{\lvert}\Pi[S(\mathrm{e}^{\mathrm{i}\theta})]-\Pi[S_{N}(\mathrm{e}^{\mathrm{i}\theta})]\big{\rvert}=\Bigg{\lvert}\sum_{n=\frac{N}{2}+1}^{\infty}a_{n}\mathrm{e}^{\mathrm{i}n\theta}\Bigg{\rvert}\leq\sum_{n=\frac{N}{2}+1}^{\infty}\lvert a_{n}\rvert\leq M(r)\sum_{n=\frac{N}{2}+1}^{\infty}r^{-n}=\frac{M(r)}{r^{N/2}(r-1)}\quad(\theta\in(-\pi,\pi]).$ (4.25)
To obtain a similar bound for the difference between $\Pi[S_{N}(\mathrm{e}^{\mathrm{i}\theta})]$ and $\Pi[\tilde{S}_{N}(\mathrm{e}^{\mathrm{i}\theta})]$, we recall the discrete Poisson summation formula [23, Chapter 6]
$\tilde{a}_{n}=a_{n}+\sum_{m\in\mathbb{Z}\setminus\{0\}}a_{n+Nm}.$ (4.26)
Together, (4.19) and (4.26) give
$\lvert\tilde{a}_{n}-a_{n}\rvert=\Bigg{\lvert}\sum_{m\in\mathbb{Z}\setminus\{0\}}a_{n+Nm}\Bigg{\rvert}\leq\sum_{m\in\mathbb{Z}\setminus\{0\}}\lvert a_{n+Nm}\rvert\leq 2M(r)r^{-n}\sum_{m=1}^{\infty}r^{-Nm}=\frac{2M(r)}{r^{n}(r^{N}-1)}\quad(\lvert n\rvert\leq N).$ (4.27)
It then follows that
$\big{\lvert}\Pi[S_{N}(\mathrm{e}^{\mathrm{i}\theta})]-\Pi[\tilde{S}_{N}(\mathrm{e}^{\mathrm{i}\theta})]\big{\rvert}=$ $\;\Bigg{\lvert}\frac{1}{2}(a_{0}-\tilde{a}_{0})+\sum_{n=1}^{N}(a_{n}-\tilde{a}_{n})\mathrm{e}^{\mathrm{i}n\theta}\Bigg{\rvert}\leq\frac{1}{2}\lvert a_{0}-\tilde{a}_{0}\rvert+\Bigg{\lvert}\sum_{n=1}^{N}(a_{n}-\tilde{a}_{n})z^{n}\Bigg{\rvert}$
$\leq$ $\;\frac{1}{2}\lvert a_{0}-\tilde{a}_{0}\rvert+\sum_{n=1}^{N}\lvert a_{n}-\tilde{a}_{n}\rvert\leq\frac{M(r)}{r^{N}-1}\Bigg{(}1+2\sum_{n=1}^{N}r^{-n}\Bigg{)}$
$=$ $\;\frac{M(r)}{r^{N}-1}\bigg{(}1+2\frac{r^{N}-1}{r^{N}(r-1)}\bigg{)}=M(r)\frac{r^{N+1}+r^{N}-2}{r^{N}(r^{N}-1)(r-1)}<M(r)\frac{r+2}{r^{N}(r-1)}\quad(\theta\in(-\pi,\pi]),$ (4.28)
where we have used that $(r^{N+1}-1)/(r^{N}-1)<r+1$ in the final step. Hence, (4.25) and (4.3) and the triangle inequality imply
$\big{\lvert}\Pi[S(\mathrm{e}^{\mathrm{i}\theta})]-\Pi[\tilde{S}_{N}(\mathrm{e}^{\mathrm{i}\theta})]\big{\rvert}\leq M(r)\frac{r^{N/2}+r+2}{r^{N}(r-1)}<\frac{4M(r)}{r^{N/2}(r-1)}\quad(\theta\in(-\pi,\pi]).$ (4.29)
We can now compute
$\big{\lvert}Q(\mathrm{e}^{\mathrm{i}\theta})-\tilde{Q}_{1,N}(\mathrm{e}^{\mathrm{i}\theta})\big{\rvert}=$ $\;\big{\lvert}\exp\big{(}\Pi[S(\mathrm{e}^{\mathrm{i}\theta})]\big{)}-\exp\big{(}\Pi[\tilde{S}_{N}(\mathrm{e}^{\mathrm{i}\theta})]\big{)}\big{\rvert}$
$=$ $\;\big{\lvert}\exp\big{(}\Pi[S(\mathrm{e}^{\mathrm{i}\theta})]\big{)}\big{(}1-\exp\big{(}\Pi[\tilde{S}_{N}(\mathrm{e}^{\mathrm{i}\theta})]-\Pi[{S}(\mathrm{e}^{\mathrm{i}\theta})]\big{)}\big{\rvert}$
$<$ $\;\exp\bigg{(}\frac{4M(r)}{r^{N/2}(r-1)}\bigg{)}-1\quad(\theta\in(-\pi,\pi]),$ (4.30)
where we have used $\lvert Q(\mathrm{e}^{\mathrm{i}\theta})\rvert=\big{\lvert}\exp\big{(}\Pi[S(\mathrm{e}^{\mathrm{i}\theta})]\big{)}\big{\rvert}<1$ and (4.29) in the final step.
Next, we use (3.11), (3.12), and (4.3) to write
$\lvert q_{n}-\tilde{q}_{n}\rvert=\frac{1}{N}\Bigg{\lvert}\sum_{m=0}^{N-1}\big{(}Q(\omega_{N}^{m})-\tilde{Q}_{1,N}(\omega_{N}^{m})\big{)}\omega_{N}^{-nm}\Bigg{\rvert}<\exp\bigg{(}\frac{4M(r)}{r^{N/2}(r-1)}\bigg{)}-1\quad(n\in[d]_{0}).$ (4.31)
We are guaranteed that the argument of the exponential in (4.31) is upper-bounded by unity provided that
$N\geq\frac{2}{\log r}\log\bigg{(}\frac{4M(r)}{r-1}\bigg{)}.$ (4.32)
Suppose that (4.32) holds. Then, using $\mathrm{e}^{x}-1<2x$ for $x\in(0,1)$ we have
$\exp\bigg{(}\frac{4M(r)}{r^{N/2}(r-1)}\bigg{)}-1<\frac{8M(r)}{r^{N/2}(r-1)}.$ (4.33)
Putting (4.33) in (4.31) yields
$\lvert q_{n}-\tilde{q}_{n}\rvert<8M(r)\frac{1}{r^{N/2}(r-1)}\quad(n\in[d]_{0})$ (4.34)
and we find that (4.34) is upper-bounded by $\varepsilon$ provided that
$N\geq\frac{2}{\log r}\log\bigg{(}\frac{8M(r)}{r-1}\frac{1}{\varepsilon}\bigg{)},$ (4.35)
which implies (4.32), holds. We can make (4.35) more precise by specifying an $r\in(1,R)$ and bounding $M(r)$. We choose $r=r_{\delta}$, defined in (4.9). We have $r_{\delta}>1$ by the assumption that $\delta\in(0,1)$. The following lemma shows that $r_{\delta}<R$ by bounding $\lvert P(z)P^{*}(1/z)\rvert$ in the annulus $A(r_{\delta})$ (4.20).
Lemma 4.3.2.
The following inequality holds,
$\lvert P(z)P^{*}(1/z)\rvert\leq 1-\delta\quad(z\in A(r_{\delta})).$ (4.36)
Proof.
We define the reciprocal polynomial to $P$ by
${P}^{\mathrm{R}}(z)\coloneqq z^{d}P^{*}(1/z);$ (4.37)
${P}^{\mathrm{R}}$ is a polynomial and hence entire. Note that $\lVert{P}^{\mathrm{R}}(z)\rVert_{\infty,\mathbb{T}}=\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 1-\delta$. It follows, by the maximum modulus principle, that
$\lvert P(z)\rvert,\lvert{P}^{\mathrm{R}}(z)\rvert\leq 1-\delta\quad(z\in\mathbb{D}).$ (4.38)
Using the conformal map $z\mapsto 1/z$, we deduce from (4.38) the corresponding bounds
$\lvert P(1/z)\rvert,\lvert{P}^{\mathrm{R}}(1/z)\rvert\leq 1-\delta\quad(z\in\mathbb{C}\setminus\overline{\mathbb{D}}).$ (4.39)
Because $P(z)=z^{d}({P}^{\mathrm{R}})^{*}(1/z)$, we have
$\lvert P(z)\rvert=\lvert z\rvert^{d}\lvert{P}^{\mathrm{R}}(1/z)\rvert\leq(1-\delta)\lvert z\rvert^{d}\quad(z\in\mathbb{C}\setminus\overline{\mathbb{D}}),$ (4.40)
where we have used (4.39). Again using $z\mapsto 1/z$, we have
$\lvert P(1/z)\rvert\leq(1-\delta)\lvert z\rvert^{-d}\quad(z\in\mathbb{D}).$ (4.41)
Combining (4.38)–(4.39) and (4.40)–(4.41) yields
$\lvert P(z)P^{*}(1/z)\rvert\leq(1-\delta)^{2}\begin{cases}\lvert z\rvert^{-d}&z\in\mathbb{D}\\ \lvert z\rvert^{d}&z\in\mathbb{C}\setminus\overline{\mathbb{D}}.\end{cases}$ (4.42)
Within the annulus $A(r_{\delta})$, this implies
$|P(z)P^{*}(1/z)|\leq(1-\delta)^{2}\frac{1}{1-\delta}=1-\delta,$ (4.43)
as desired. ∎
The next lemma provides an estimate for $M(r_{\delta})$ (4.18).
Lemma 4.3.3.
The following bound holds,
$M(r_{\delta})\leq\log\big{(}\tfrac{1}{\delta}\big{)}.$ (4.44)
Proof.
We write
$M(r_{\delta})=\max_{z\in\partial A(r_{\delta})}\big{\lvert}\log\big{(}1-P(z)P^{*}(1/z)\big{)}\big{\rvert}.$ (4.45)
Due the estimate (4.36), which guarantees $\lvert P(z)P^{*}(1/z)\rvert\leq 1-\delta<1$ on $\overline{A(r_{\delta})}$, we may use the Maclaurin series for $\log(1-z)$ and (4.36) to write
$\big{\lvert}\log\big{(}1-P(z)P^{*}(1/z)\big{)}\big{\rvert}=$ $\;\Bigg{\lvert}\sum_{n=1}^{\infty}\frac{(P(z)P^{*}(1/z))^{n}}{n}\Bigg{\rvert}\leq\sum_{n=1}^{\infty}\frac{\lvert P(z)P^{*}(1/z)\rvert^{n}}{n}$
$\leq$ $\;\sum_{n=1}^{\infty}\frac{(1-\delta)^{n}}{n}=\log\big{(}\tfrac{1}{\delta}\big{)}\quad(z\in\overline{A(r_{\delta})});$ (4.46)
the result (4.44) follows. ∎
Putting (4.9) and (4.44) into (4.35) gives the result (4.8).
Next, we analyze the asymptotic behaviour of our bound (4.8) on a sufficient $N$. For this we first introduce the parameter
$c_{\delta}\coloneqq\frac{1}{\log\frac{1}{1-\delta}}$ (4.47)
and compute
$N_{0}(\varepsilon,\delta,d)=\bigg{\lceil}2dc_{\delta}\log\frac{1}{\varepsilon}+2dc_{\delta}\log\left(8\log\frac{1}{\delta}\right)+2dc_{\delta}\log\left(\frac{1}{r_{\delta}-1}\right)\bigg{\rceil}$ (4.48)
Immediately, the asymptotic
$N_{0}(\varepsilon,\delta,d)\sim 2dc_{\delta}\log\frac{1}{\varepsilon}=O\!\left(\log\frac{1}{\varepsilon}\right),\quad(\varepsilon\to 0,\delta,d\ \text{fixed})$ (4.49)
follows; this regime describes increasing the accuracy of the algorithm for a fixed polynomial.
Next, note that $c_{\delta}d\to+\infty$ as $\delta\to 0^{+}$ and/or $d\to\infty$, such that we have the asymptotic equation
$\frac{1}{r_{\delta}-1}=\frac{1}{\exp(\frac{1}{c_{\delta}d})-1}\sim c_{\delta}d\quad(\delta\to 0\ \text{and/or}\ d\to\infty).$ (4.50)
We insert this into (4.48), but keep $\varepsilon$ because we will take a limit $\varepsilon\to 0$ later:
$N_{0}(\varepsilon,\delta,d)$ $\sim 2dc_{\delta}\log\frac{1}{\varepsilon}+2dc_{\delta}\log\left(8\log\frac{1}{\delta}\right)+2dc_{\delta}\log\left(c_{\delta}d\right)\quad(\delta\to 0\ \text{and/or}\ d\to\infty).$ (4.51)
The middle term is subdominant: In the case $d\to\infty$, the middle logarithm is a constant, and in the case $\delta\to 0$ note that $c_{\delta}\sim 1/\delta$. Dropping the middle term results in
$N_{0}(\varepsilon,\delta,d)\sim 2dc_{\delta}\log\left(\frac{c_{\delta}d}{\varepsilon}\right)\quad(\delta\to 0\ \text{and/or}\ d\to\infty).$ (4.52)
For fixed $\delta$, we retrieve the joint asymptotic
$N_{0}(\varepsilon,\delta,d)\sim 2dc_{\delta}\log\frac{d}{\varepsilon}=O\!\left(d\log\frac{d}{\varepsilon}\right)\quad(\varepsilon\to 0,d\to\infty,\delta\ \text{fixed}).$ (4.53)
As $\delta\to 0$, we get
$N_{0}(\varepsilon,\delta,d)\sim 2\frac{d}{\delta}\log\left(\frac{d}{\varepsilon\delta}\right)\quad(\delta\to 0).$ (4.54)
which is valid regardless of whether $\varepsilon\to 0$ and/or $d\to\infty$. In particular, (4.11) holds provided at least one of $\varepsilon\to 0$, $\delta\to 0$, or $d\to\infty$.
4.4 Proof of Corollary 3.1
Let $Q$ be the exact complementary polynomial to $P$ obtained from Corollary 2.2 in the form (3.9).
We first prove (4.12). From (4.10), we have
$\lVert Q(z)-\tilde{Q}_{2,N}(z)\rVert_{\mathbb{T},\infty}<(d+1)\varepsilon$ (4.55)
and hence,
$\big{\lVert}|Q(z)|^{2}-|\tilde{Q}_{2,N}(z)|^{2}\big{\rVert}_{\infty,\mathbb{T}}=$ $\;\lVert 2Q^{\prime}(z)-Q^{\prime}(z)+\tilde{Q}(z)\rVert_{\infty,\mathbb{T}}\cdot\big{\lVert}Q(z)-\tilde{Q}_{2,N}(z)\big{\rVert}_{\infty,\mathbb{T}}$
$<$ $\;(2+(d+1)\epsilon)(d+1)\varepsilon)<(d+1)(d+3)\varepsilon.$ (4.56)
$\Phi(P,\tilde{Q}_{2,N})=$ $\;\big{\lVert}\lvert P(z)\rvert^{2}+\lvert\tilde{Q}_{2,N}(z)\rvert^{2}-1\big{\rVert}_{\infty,\mathbb{T}}=\big{\lVert}\big{(}\lvert P(z)\rvert^{2}+\lvert Q(z)\rvert^{2}-1\big{)}+\big{(}\lvert\tilde{Q}_{2,N}(z)\rvert^{2}-\lvert Q(z)\rvert^{2}\big{)}\big{\rVert}_{\infty,\mathbb{T}}$
$\leq$ $\;\Phi(P,Q)+\big{\lVert}\lvert\tilde{Q}_{2,N}(z)\rvert^{2}-\lvert Q(z)\rvert^{2}\big{\rVert}_{\infty,\mathbb{T}}=\big{\lVert}\lvert\tilde{Q}_{2,N}(z)\rvert^{2}-\lvert Q(z)\rvert^{2}\big{\rVert}_{\infty,\mathbb{T}}<(d+1)(d+3)\varepsilon,$ (4.57)
which is (4.12).
We next prove (4.13). Inserting
$\tilde{q}_{n+m}\tilde{q}_{m}^{*}=$ $\;\big{(}{q}_{n+m}+(\tilde{q}_{n+m}-q_{n+m})\big{)}\big{(}q_{m}^{*}+(\tilde{q}_{m}^{*}-q_{m}^{*})\big{)}$
$=$ $\;q_{n+m}q_{m}^{*}+q_{n+m}(\tilde{q}_{m}^{*}-q_{m}^{*})+q_{m}^{*}(\tilde{q}_{n+m}-q_{n+m})+(\tilde{q}_{n+m}-q_{n+m})(\tilde{q}_{m}^{*}-q_{m}^{*});$ (4.58)
into the summand of the loss function (4.2) gives
$p_{n+m}p_{m}^{*}+\tilde{q}_{n+m}\tilde{q}_{m}^{*}-{\delta_{n,0}}=\\ \big{(}p_{n+m}p_{m}^{*}+{q}_{n+m}{q}_{m}^{*}-{\delta_{n,0}}\big{)}+q_{n+m}(\tilde{q}_{m}^{*}-q_{m}^{*})+q_{m}^{*}(\tilde{q}_{n+m}-q_{n+m})+(\tilde{q}_{n+m}-q_{n+m})(\tilde{q}_{m}^{*}-q_{m}^{*}).$ (4.59)
By the $\ell^{1}$$\ell^{2}$ norm inequality, we have
$\tilde{\Phi}(P,\tilde{Q}_{2,N})\leq\sum_{n=-d}^{d}\Bigg{\lvert}\sum_{m=0}^{d}\big{(}p_{n+m}p_{m}^{*}+\tilde{q}_{n+m}\tilde{q}_{m}^{*}-{\delta_{n,0}}\big{)}\Bigg{\rvert}.$ (4.60)
Using (4.59), it follows that
$\tilde{\Phi}(P,\tilde{Q}_{2,N})\leq$ $\;\sum_{n=-d}^{d}\Bigg{\lvert}\sum_{m=0}^{d}\big{(}p_{n+m}p_{m}^{*}+{q}_{n+m}{q}_{m}^{*}-{\delta_{n,0}}\big{)}\Bigg{\rvert}+\sum_{n=-d}^{d}\Bigg{\lvert}\sum_{m=0}^{d}q_{n+m}(\tilde{q}_{m}^{*}-q_{m}^{*})\Bigg{\rvert}$
$\;+\sum_{n=-d}^{d}\Bigg{\lvert}\sum_{m=0}^{d}q_{m}^{*}(\tilde{q}_{n+m}-q_{n+m})\Bigg{\rvert}+\sum_{n=-d}^{d}\Bigg{\lvert}\sum_{m=0}^{d}(\tilde{q}_{n+m}-q_{n+m})(\tilde{q}_{m}^{*}-q_{m}^{*})\Bigg{\rvert}$ (4.61)
The first term in (4.4) is seen to be zero by again appealing to the $\ell^{1}$$\ell^{2}$ norm inequality and using the fact that $\Phi(P,Q)=0$. The remaining terms are bounded as follows. By writing
$q_{n}=\frac{1}{2\pi\mathrm{i}}\int_{\mathbb{T}}\Bigg{(}\sum_{m=0}^{d}q_{m}z^{m}\Bigg{)}\frac{\mathrm{d}z}{z^{n+1}}\quad(n\in[d]_{0}),$ (4.62)
we see that
$\lvert q_{n}\rvert\leq\lVert Q(z)\rVert_{\infty,\mathbb{T}}\leq 1\quad(n\in[d]_{0}).$ (4.63)
Hence, (4.4) with (4.63) and the bound on $|q_{n}-\tilde{q}_{n}|$ from Theorem 3 gives
$\tilde{\Phi}(P,\tilde{Q}_{2,N})\leq(2d+1)(d+1)\varepsilon+(2d+1)(d+1)\varepsilon+(2d+1)(d+1)\varepsilon^{2}<3(2d+1)(d+1)\varepsilon,$ (4.64)
which is (4.13).
4.5 Proof of Theorem 4
First, note that we have
$\big{\lVert}|P(z)|^{2}-|(1-\tfrac{\varepsilon}{4})P(z)|^{2}\big{\rVert}_{\infty,\mathbb{T}}\leq\big{\lVert}P(z)+(1-\tfrac{\varepsilon}{4})P(z)\big{\rVert}_{\infty,\mathbb{T}}\cdot\big{\lvert}P(z)-(1-\tfrac{\varepsilon}{4})P(z)\big{\rVert}_{\infty,\mathbb{T}}<\frac{\varepsilon}{2}.$ (4.65)
Let $Q$ be the exact canonical complementary polynomial to $(1-\tfrac{\varepsilon}{4})P$ in the form (3.9). From Theorem 3, it follows that
$|q_{n}-\tilde{q}_{n}|<\frac{\varepsilon}{5(d+1)}\quad(n\in[d]_{0}),$ (4.66)
which implies that
$\lVert Q(z)-\tilde{Q}_{2,N}(z)\rVert_{\infty,\mathbb{T}}<\frac{\varepsilon}{5}$ (4.67)
and
$\big{\lVert}|Q(z)|^{2}-|\tilde{Q}_{2,N}(z)|^{2}\big{\rVert}_{\infty,\mathbb{T}}\leq\lVert 2Q(z)-Q(z)+\tilde{Q}_{2,N}(z)\rVert_{\infty,\mathbb{T}}\cdot\lVert Q(z)-\tilde{Q}_{2,N}(z)\rVert_{\infty,\mathbb{T}}\leq\bigg{(}2+\frac{\varepsilon}{5}\bigg{)}\frac{\varepsilon}{5}<\frac{\varepsilon}{2}.$ (4.68)
Using (4.65), (4.68), and the fact that $\Phi((1-\tfrac{\varepsilon}{4})P,Q)=0$, we obtain
$\Phi(P,\tilde{Q}_{2,N})=\big{\lVert}|P(z)|^{2}+|\tilde{Q}_{2,N}(z)|^{2}-1\big{\rVert}_{\infty,\mathbb{T}}<\frac{\varepsilon}{2}+\frac{\varepsilon}{2}+\big{\lVert}|(1-\tfrac{\varepsilon}{4})P(z)|^{2}+|Q(z)|^{2}-1\big{\rVert}_{\infty,\mathbb{T}}=\varepsilon.$ (4.69)
The scaling (4.16) can be obtained by inserting $\varepsilon\to\varepsilon/5(d+1)$ and $\delta\to\varepsilon/4$ into (4.11).
5 Numerical results
In practice, our new algorithm for computation of complementary polynomials is extremely fast and accurate. A reference implementation of Algorithm 1 in Python using the FFT from the PyTorch library [24] is provided in Figure 1. While our proofs in the previous section focus on $N$ even, we run the algorithm for both odd and even choices of $N$; our results carry over. All benchmarks are performed on an M2 Macbook Pro with 16 GB RAM; we have not used any GPU acceleration that PyTorch optionally provides for the FFT.
To facilitate comparison with prior work, we compute the same loss function (4.2) as used in the optimization based method [3]. Here, we just compute the loss function $\tilde{\Phi}$ of the algorithm’s output as a measure of how well it satisfies the complementarity condition, see Proposition 4.1.1 and Corollary 3.1. In [3], the coefficients of an approximate complementary polynomial are determined by an optimization procedure minimizing the loss function.
Figure 2(a) demonstrates Algorithm 1 for random polynomials with $\delta=0.2$, i.e. $\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 0.8$. We consider random polynomials up to degree $d=10^{7}$ and can numerically confirm the asymptotic scalings $N=O\left(d\log(1/\varepsilon)\right)$ from Theorem 3, up to logarithmic terms in $d$. It appears that the bound on $N$ in Theorem 3 is not tight and lower $N$ are sufficient in practice. Already at $N=4d$, our algorithm achieves a far better loss function than the optimization based method [3]. The runtime of Algorithm 1 is shown in Figure 3. We observe a runtime $O(N\log N)$ as expected from the FFTs. Comparing to [3]’s algorithm running on a CPU, we achieve far better runtimes for the same value of the loss function.
Beyond its numerical efficacy, Algorithm 1 provably and reproducibly targets the same canonical solution, where $Q$ has no roots in $\mathbb{\mathbb{D}}$. This is in contrast to the optimization based algorithm from [3], which can (and does) converge to very different solutions $Q(z)$ for the same target polynomial $P(z)$ on successive runs. There is also no guarantee that the optimization doesn’t find a local minimum.
For polynomials with $\delta=0$, $Q_{0}(z)\neq 1$ because $1-|P(z)|^{2}$ has roots on $z\in\mathbb{T}$. Theorem 4 proves that in those cases, Algorithm 2 provides a solution, scaling as $N=O\!\left(\frac{d}{\varepsilon}\log\frac{d}{\varepsilon}\right)$ for desired error $\varepsilon$. Instead of running Algorithm 2 including initial downscaling of the polynomial, here we simply run Algorithm 1 despite lack of rigorous justification and show the results in Figure 2(b). We are running the same code of the reference implementation (Figure 1) used for $\delta=0.2$. Even though Algorithm 1 assumes $Q_{0}(z)=1$, which is not the case any more, it still seems to give good results, even at degree $d=10^{7}$. In fact, we can still achieve losses better than those of [3] at significantly lower runtimes. The computation of $\log(1-|P(\omega_{N}^{n})|^{2})$ in step 2 of Algorithm 1 does not cause any issues, despite $1-|P(z)|^{2}$ having roots on $z\in\mathbb{T}$, because, generically, these will not be located exactly at roots of unity $\omega_{N}^{n}$ (otherwise, the polynomial can be rotated $P(z)\to P(\mathrm{e}^{i\alpha}z)$ by a suitable phase $\mathrm{e}^{i\alpha}$). Empirically we find a scaling of the algorithm as $N\approx O(1/\sqrt[4]{\varepsilon^{\prime}})$ for a desired accuracy $\varepsilon^{\prime}$ in the loss function $\tilde{\Phi}$; proving this could be the subject of future work.
import torch
def complementary(poly, N):
    """Algorithm 1 to compute the complementary polynomial
    Parameters:
    poly : length (d+1) vector of monomial coefficients of P(z)
    N int : size of the FFT, N >= (d+1)
    Returns:
    length (d+1) vector of monomial coefficients of Q(z)"""
    # Pad P to FFT dimension N
    paddedPoly = torch.zeros(N, dtype=torch.complex128)
    paddedPoly[:poly.shape[0]] = poly
    # Evaluate P(omega) at roots of unity omega
    pEval = torch.fft.ifft(paddedPoly, norm="forward")
    # Compute log(1-|P(omega)|^2) at roots of unity omega
    theLog = torch.log(1-torch.square(torch.abs(pEval)))
    # Apply Fourier multiplier in Fourier space
    modes = torch.fft.fft(theLog, norm="forward")
    modes[0] *= 1/2 # Note modes are ordered differently in the text
    modes[N//2+1:] = 0
    theLog = torch.fft.ifft(modes, norm="forward")
    # Compute coefficients of Q
    coefs = torch.fft.fft(torch.exp(theLog), norm="forward")
    # Truncate to length of Q polynomial
    q = coefs[:poly.shape[0]]
    return q
Figure 1: Reference implementation in Python of Algorithm 1 for finding the complementary polynomial. Input and output are (complex) coefficient vectors of $P(z)$ and $Q(z)$ in the monomial bases, along with integer $N$ controlling the accuracy of the output.
(a) Polynomials with $\delta=0.2$. We observe the approximate scaling $N\approx O(d\log 1/\tilde{\Phi})$, linear scaling with the degree $d$ and logarithmic scaling with the accuracy $\tilde{\Phi}$. This is commensurate with our proof of error scaling (Theorem 3) for Algorithm 1, up to logarithmic factors in $d$.
(b) Polynomials with $\delta=0$. Despite lack of rigorous justification for Algorithm 1 in this regime (necessitating the use of Algorithm 2, slightly scaling down the polynomial), we observe that Algorithm 1 is efficacious with a scaling behaviour of $N\approx O(d/\sqrt[4]{\tilde{\Phi}})$ in the accuracy $\tilde{\Phi}$.
Figure 2: Finding the complementary polynomial with the reference implementation (Figure 1) of Algorithm 1 for random polynomials of various degrees $d$. We plot the achieved loss function (4.2) by chosen FFT dimension $N$. The optimization code from GQSP [3] only reaches losses up to $10^{-6}$, which Algorithm 1 achieves at far faster runtimes (see Figure 3). In the plots, loss functions saturate around $10^{-14}$ due to double-precision floating-point arithmetics.
Figure 3: Runtime of the reference implementation (Figure 1) of Algorithm 1 on an M2 Macbook Pro with 16GB RAM. We observe the scaling $N\log N$ due to the fast Fourier transformations.
6 Discussion
In this paper, we have addressed the solvability of the complementary polynomials problem, Problem 1. Our main mathematical result, Theorem 2, is an exact representation, written as a set of contour integrals, for the complementary polynomial throughout the entire complex plane. We use a Fourier analytic variant of Theorem 2, Corollary 2.2, as a basis for developing a numerical method to obtain the complementary polynomial explicitly in the monomial basis.
We give the following closing remarks.
  1. 1.
    Problem 1 is a special case of the Fejér-Riesz problem. Our methods for solving Problem 1 both analytically and numerically are equally applicable to the more general Féjer-Riesz problem.
  2. 2.
    We constructed integral representations of $Q$ on the entire complex plane. It is interesting to consider if these integrals might be explicitly computable. It has been shown [25] that the real line Hilbert transforms of logarithms of polynomials are expressible in terms of the roots of the polynomial. As the integration in (1.4b) amounts to a periodic Hilbert transformation of a logarithm of a Laurent polynomial, we have obtained an analogous result by comparing (1.4b) with (2.1). As any root finding algorithm is anticipated to be more expensive than Algorithm 1, this observation is not of practical consequence.
  3. 3.
    While we have chosen the representation of $Q$ on $\mathbb{T}$ (1.4b) and its Fourier analytic equivalent (1.6) as the basis for our numerical method, it is also possible to construct $Q$ though interpolation at any $d+1$ distinct points in the complex plane. Theorem 2 provides a basis to do this. The representations (1.4a) and (1.4c) are essentially Cauchy transforms, which the results of [26] show are computable $O(N\log N)$ time. However, the interpolation through $z$ values that are not roots of unity will be more expensive as the FFT is not applicable.
  4. 4.
    There are interesting connections between the present work and the classical signal processing literature. In the context of pulse synthesis, generically non-polynomial solutions to (1.1) have been investigated on both $\mathbb{R}$ [27] and $\mathbb{T}$ [28]. We particularly highlight that in [28], a formula similar to (1.4b) appears, apparently obtained by conformal mapping of an analogous formula on the real line in [27]. In the context of phase retrieval, similar technologies as in this paper, namely Hilbert transforms of logarithms, have been employed [29].
  5. 5.
    In [3], the complementary polynomials problem was rephrased as an optimization problem over the coefficients of $Q$. The corresponding numerical method was based on minimizing the objective function (4.2). Equation (4.2) defines a system of nonlinear algebraic equations for the coefficients $(q_{n})_{n=0}^{d}$ and it is an interesting question whether this system could be solved explicitly using discrete mathematics or finite-dimensional linear algebra, without appealing to analysis as in this paper. While we make no claim this is impossible, we note that Problem 1 is essentially a special case of the Fejér-Riesz problem, for which we are unaware of an explicit solution by such means.
  6. 6.
    In the proof of Theorem 2, we have characterized the distinct solutions of Problem 1 and isolated the canonical solution having all roots outside the unit disk. The fact that the algorithm proposed in [3] does not target a particular solution may explain, in part, its effectiveness. It is an interesting question to consider the effect of the particular solution targeted on the resulting phase factors associated to a pair of complementary polynomials $(P,Q)$. In particular, does the geometry of the roots of $Q$ affect the structure of the phase factors?
Acknowledgements
We thank Earl Campbell, György Geher, Tharon Holdsworth, Tanuj Khattar, Nicholas Rubin, and especially Hari Krovi for useful discussions. This work was partially funded by Innovate UK (grant reference 10071684).
Appendix A QSP conventions
In this appendix, we substantiate our claim that constructing complementary polynomials in different QSP conventions may be achieved by using the results of Theorem 2 in conjunction with mappings between the conventions. We first show, in Section A.1, that GQSP implies the Laurent formulation of QSP due to Haah [11]. In Section A.2, we show that the Laurent formulation of QSP implies the standard formulation of QSP; see [6] for a similar presentation of interrelations between different parameterizations of standard QSP.
A.1 From GQSP to Laurent QSP
We begin by computing the remaining entries in the target matrix of GQSP.
Proposition A.1.1.
The precise form of the matrix realized by the GQSP sequence (1.2) is
$\left(\begin{array}[]{cc}P(z)&Q(z)\\ &*\end{array}\right)=\left(\begin{array}[]{cc}P(z)&Q(z)\\ u_{d}(z)Q^{*}(1/z)&-u_{d}(z)P^{*}(1/z)\end{array}\right),$ (A.1)
where
$u_{d}(z)=(-1)^{d}z^{d}\mathrm{e}^{\mathrm{i}\big{(}\lambda+\sum_{j=0}^{d}\phi_{j}\big{)}}.$ (A.2)
Proof.
By unitarity, (A.1) must hold for some function $u_{d}:\mathbb{T}\to\mathbb{T}$. Using (1.1), (A.1), and
$\det\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}(\lambda+\phi_{0})}\cos\theta_{0}&\mathrm{e}^{\mathrm{i}\lambda}\sin\theta_{0}\\ \mathrm{e}^{\mathrm{i}\theta_{0}}\sin\theta_{0}&-\cos\theta_{0}\end{array}\right)=-\mathrm{e}^{\mathrm{i}(\lambda+\phi_{0})},\quad\det\left(\begin{array}[]{cc}z&0\\ 0&1\end{array}\right)=z,\quad\det\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}\phi_{j}}\cos\theta_{j}&\sin\theta_{j}\\ \mathrm{e}^{\mathrm{i}\phi_{j}}\sin\theta_{j}&-\cos\theta_{j}\end{array}\right)=-\mathrm{e}^{\mathrm{i}\phi_{j}}$ (A.3)
to evaluate the determinant of both sides of (1.2) gives (A.2). ∎
By combining Theorem 1 and Proposition A.1.1 and making the replacement $z\to z^{2}$, we have
$\left(\begin{array}[]{cc}P(z^{2})&Q(z^{2})\\ u_{d}(z^{2})Q^{*}(1/z^{2})&-u_{d}(z^{2})P^{*}(1/z^{2})\end{array}\right)=\\ \left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}(\lambda+\phi_{0})}\cos\theta_{0}&\mathrm{e}^{\mathrm{i}\lambda}\sin\theta_{0}\\ \mathrm{e}^{\mathrm{i}\phi_{0}}\sin\theta_{0}&-\cos\theta_{0}\end{array}\right)\Bigg{[}\prod_{j=1}^{d}\left(\begin{array}[]{cc}z^{2}&0\\ 0&1\end{array}\right)\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}\phi_{j}}\cos\theta_{j}&\sin\theta_{j}\\ \mathrm{e}^{\mathrm{i}\phi_{j}}\sin\theta_{j}&-\cos\theta_{j}\end{array}\right)\Bigg{]}\quad(z\in\mathbb{T}).$ (A.4)
We rename parameters $\lambda\to\lambda_{0}$ and $\phi_{j}\to\phi_{j}+\lambda_{j+1}$ for $j\in[d-1]_{0}$ in (A.4). This yields
$\left(\begin{array}[]{cc}P(z^{2})&Q(z^{2})\\ u_{d}(z^{2})Q^{*}(1/z^{2})&-u_{d}(z^{2})P^{*}(1/z^{2})\end{array}\right)=\\ \left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}(\lambda_{0}+\phi_{0}+\lambda_{1})}\cos\theta_{0}&\mathrm{e}^{\mathrm{i}\lambda_{0}}\sin\theta_{0}\\ \mathrm{e}^{\mathrm{i}\phi_{0}+\lambda_{1}}\sin\theta_{0}&-\cos\theta_{0}\end{array}\right)\Bigg{[}\prod_{j=1}^{d-1}\left(\begin{array}[]{cc}z^{2}&0\\ 0&1\end{array}\right)\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}(\phi_{j}+\lambda_{j+1})}\cos\theta_{j}&\sin\theta_{j}\\ \mathrm{e}^{\mathrm{i}(\phi_{j}+\lambda_{j+1})}\sin\theta_{j}&-\cos\theta_{j}\end{array}\right)\Bigg{]}\\ \times\left(\begin{array}[]{cc}z^{2}&0\\ 0&1\end{array}\right)\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}\phi_{d}}\cos\theta_{j}&\sin\theta_{j}\\ \mathrm{e}^{\mathrm{i}\phi_{d}}\sin\theta_{j}&-\cos\theta_{j}\end{array}\right).$ (A.5)
Using the factorizations
$\begin{split}\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}(\lambda_{0}+\phi_{0}+\lambda_{1})}\cos\theta_{0}&\mathrm{e}^{\mathrm{i}\lambda_{0}}\sin\theta_{0}\\ \mathrm{e}^{\mathrm{i}(\phi_{0}+\lambda_{1})}\sin\theta_{0}&-\cos\theta_{0}\end{array}\right)=&\;\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}(\lambda_{0}+\phi_{0})}\cos\theta_{0}&\mathrm{e}^{\mathrm{i}\lambda_{0}}\sin\theta_{0}\\ \mathrm{e}^{\mathrm{i}\phi_{0}}\sin\theta_{0}&-\cos\theta_{0}\end{array}\right)\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}\lambda_{1}}&0\\ 0&1\end{array}\right),\\ \left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}(\phi_{j}+\lambda_{j+1})}\cos\theta_{j}&\sin\theta_{j}\\ \mathrm{e}^{\mathrm{i}(\phi_{j}+\lambda_{j+1})}\sin\theta_{j}&-\cos\theta_{j}\end{array}\right)=&\;\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}\phi_{j}}\cos\theta_{j}&\sin\theta_{j}\\ \mathrm{e}^{\mathrm{i}\phi_{j}}\sin\theta_{j}&-\cos\theta_{j}\end{array}\right)\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}\lambda_{j+1}}&0\\ 0&1\end{array}\right)\end{split}$ (A.6)
and commutativity of diagonal matrices in (A.4), we deduce
$\left(\begin{array}[]{cc}P(z^{2})&Q(z^{2})\\ u_{d}(z^{2})Q^{*}(1/z^{2})&-u_{d}(z^{2})P^{*}(1/z^{2})\end{array}\right)=\\ \left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}(\lambda_{0}+\phi_{0})}\cos\theta_{0}&\mathrm{e}^{\mathrm{i}\lambda_{0}}\sin\theta_{0}\\ \mathrm{e}^{\mathrm{i}\phi_{0}}\sin\theta_{0}&-\cos\theta_{0}\end{array}\right)\Bigg{[}\prod_{j=1}^{d}\left(\begin{array}[]{cc}z^{2}&0\\ 0&1\end{array}\right)\left(\begin{array}[]{cc}\mathrm{e}^{\mathrm{i}(\lambda_{j}+\phi_{j})}\cos\theta_{j}&\mathrm{e}^{\mathrm{i}\lambda_{j}}\sin\theta_{j}\\ \mathrm{e}^{\mathrm{i}\phi_{j}}\sin\theta_{j}&-\cos\theta_{j}\end{array}\right)\Bigg{]}.$ (A.7)
We make the following specializations in (A.2) and (A.7),
$\lambda_{j}=\frac{\pi}{2}\quad\phi_{j}=\frac{\pi}{2},\quad\theta_{j}\to\theta_{j}+\pi\quad(j\in[d]_{0})$ (A.8)
and multiply (A.5) by $z^{-d}$ to obtain
$\left(\begin{array}[]{cc}z^{-d}P(z^{2})&z^{-d}Q(z^{2})\\ -z^{d}Q^{*}(1/z^{2})&z^{d}P^{*}(1/z^{2})\end{array}\right)=\\ \left(\begin{array}[]{cc}\cos\theta_{0}&\mathrm{i}\sin\theta_{0}\\ \mathrm{i}\sin\theta_{0}&\cos\theta_{0}\end{array}\right)\Bigg{[}\prod_{j=1}^{d}\left(\begin{array}[]{cc}z&0\\ 0&z^{-1}\end{array}\right)\left(\begin{array}[]{cc}\cos\theta_{j}&\mathrm{i}\sin\theta_{j}\\ \mathrm{i}\sin\theta_{j}&\cos\theta_{j}\end{array}\right)\Bigg{]}.$ (A.9)
To proceed, we recall Corollary 2.1 and the $\mathrm{U}(1)$ invariance of Problem 1. Together, these guarantee that given $P\in\mathbb{R}[z]$ satisfying the conditions of Theorem 2, there exists a $Q\in\mathrm{i}\mathbb{R}[z]$ that is complementary to $P$. By setting $F(z)=z^{-d}P(z^{2})$ and $\mathrm{i}G(z)=z^{-d}Q(z^{2})$ in (A.9), we deduce the Laurent formulation of QSP:
Theorem 5 (Laurent quantum signal processing, [11]).
Let $F\in\mathbb{R}[z,z^{-1}]$ with $\deg F=d\in\mathbb{Z}_{\geq 1}$ and parity $d\bmod 2$. Then, there exists $G\in\mathbb{R}[z,z^{-1}]$ and $(\theta_{j})_{j=0}^{d}\in(-\pi,\pi]^{d+1}$ such that
$\lvert F(z)\rvert^{2}+\lvert G(z)\rvert^{2}=1\quad(z\in\mathbb{T})$ (A.10)
and
$\left(\begin{array}[]{cc}F(z)&\mathrm{i}G(z)\\ \mathrm{i}G(z^{-1})&F(z^{-1})\end{array}\right)=\left(\begin{array}[]{cc}\cos\theta_{0}&\mathrm{i}\sin\theta_{0}\\ \mathrm{i}\sin\theta_{0}&\cos{}\theta_{0}\end{array}\right)\Bigg{[}\prod_{j=1}^{d}\left(\begin{array}[]{cc}z&0\\ 0&z^{-1}\end{array}\right)\left(\begin{array}[]{cc}\cos\theta_{j}&\mathrm{i}\sin\theta_{j}\\ \mathrm{i}\sin\theta_{j}&\cos{}\theta_{j}\end{array}\right)\Bigg{]}\quad(z\in\mathbb{T})$ (A.11)
hold.
To apply the results of this paper to determine $G(z)$, set $P(z)=z^{\frac{d}{2}}F(\sqrt{z})$; this polynomial satisfies the conditions of Problem 1. If $Q$ is the canonical complementary polynomial to $P$ with purely imaginary coefficients, we see that $\mathrm{i}G(z)=z^{-d}Q(z^{2})$.
A.2 From Laurent QSP to real QSP
To recover the formulation of standard QSP based on polynomials on $[-1,1]$, we will employ a definition of the Chebyshev polynomials of the first kind,
$T_{n}(x)\coloneqq\frac{1}{2}(z^{n}+z^{-n})\quad(z\in\mathbb{T}),$ (A.12)
where $x\coloneqq\mathrm{Re}(z)$.
Suppose that $F$ satisfies $F(z^{-1})=F(z)$, i.e.,
$F(z)=f_{0}+\frac{1}{2}\sum_{n=1}^{d}f_{n}(z^{n}+z^{-n})=\sum_{n=0}^{d}f_{n}T_{n}(x),$ (A.13)
where we have used (A.12). Thus, the $(1,1)$ entry of (A.11) can be understood as a real-valued polynomial on $[-1,1]$. This observation in conjunction with Theorem 5 gives the following result.
Theorem 6 (Real quantum signal processing).
Let $p\in\mathbb{R}[x]$ with $\deg p=d\in\mathbb{Z}_{\geq 1}$ with parity $d\bmod 2$ satisfy $\lvert p(x)\rvert\leq 1$ on $[-1,1]$. Then, there exists $G\in\mathbb{R}[z,z^{-1}]$ such that
$\lvert p(x)\rvert^{2}+\lvert G(z)\rvert^{2}=1\quad(z\in\mathbb{T},\,x=\mathrm{Re}\,z)$ (A.14)
and $(\theta_{j})_{j=0}^{d}\in(-\pi,\pi]^{d+1}$ such that
$\left(\begin{array}[]{cc}p(x)&\mathrm{i}G(z)\\ \mathrm{i}G(z^{-1})&p(x)\end{array}\right)=\left(\begin{array}[]{cc}\cos\theta_{0}&\mathrm{i}\sin\theta_{0}\\ \mathrm{i}\sin\theta_{0}&\cos{}\theta_{0}\end{array}\right)\Bigg{[}\prod_{j=1}^{d}\left(\begin{array}[]{cc}z&0\\ 0&z^{-1}\end{array}\right)\left(\begin{array}[]{cc}\cos\theta_{j}&\mathrm{i}\sin\theta_{j}\\ \mathrm{i}\sin\theta_{j}&\cos{}\theta_{j}\end{array}\right)\Bigg{]}\quad(z\in\mathbb{T},\,x=\mathrm{Re}\,z)$ (A.15)
holds.
Given $p\in\mathbb{R}[x]$ satisfying the conditions of Theorem 6, compute the Chebyshev coefficients of $p$ and denote them by $(f_{n})_{n=0}^{d}$. This defines a Laurent polynomial $F$ via (A.13). The recipe for constructing $G$ below Theorem 5 is now applicable.